OpenStudy (jayda123):

I'm having trouble understand this!:( Julie is throwing a surprise birthday party for her friend. She wants to spend no more than 35.00 on decorations. At a local party store, she can buy helium filled balloons for $2.50 each and rolls of streamers $3.50 per roll. Let x represent the number of balloons purchased and y be the number of rolls of streamers. Then the equality 2.50x +3.50 <= 35.00 graphed below determines how many each type of decoration she can buy. If Julie wants 6 balloons, how many rolls of streamers can she buy without going over budget?

OpenStudy (jdoe0001):

Julie only has $35 in her pocket, to spend in a party cost for balllons is 2.5 each roll, "x" rolls will cost 2.5 * x or 2.5x cost for streamers is 3.5 each roll, "y" rolls will cost 3.5 * x, or 3.5x whatever 2.5x plus 3.5x is, has to be less than 35, or 35 exactly because if it's more, say, 37, then Julie won't be able to pay for them recall, she only has $35 in her pocket, not 36 or 37 or 50, just 35 thus \(\large 2.5x + 3.5y \le 35\)

OpenStudy (jdoe0001):

so.. hmm the balloons are in a roll.. so... "x" baloons, will be 2.5x anyhow, now, if Julie wants to only buy 6 balloons, namely x = 6 <-- recall the "x" is quantity of balloons so, one could say \(\bf 2.5({\color{brown}{ x}})+3.5y\le 35\implies 2.5({\color{brown}{ 6}})+3.5y\le 35\impliedby \textit{solve for }y\) since "y" is the rolls of streamers once you find "y", you'd know how many streamers she can afford, after getting 6 balloons

OpenStudy (jayda123):

Ohhhh okay, i understand now! Thank you so much!!

OpenStudy (jdoe0001):

yw

OpenStudy (jayda123):

So If I rounded it which was, 5.7, it would be 6